Two cubical dice are thrown and their scores addes together if X= add is the sum of the scores on the two dice what is P(X is divided by 4)

1
Expert's answer
2022-02-07T17:47:06-0500

When rolling two dice, there are "6^2=36" outcomes for scores


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 1+1 & 1+2 & 1+3 & 1+4 & 1+5 & 1+6 \\
 \hdashline
 2 & 2+1 & 2+2 & 2+3 & 2+4 & 2+5 & 2+6 \\
 \hdashline
 3 & 3+1 & 3+2 & 3+3 & 3+4 & 3+5 & 3+6 \\
 \hdashline
 4 & 4+1 & 4+2 & 4+3 & 4+4 & 4+5 & 4+6 \\
 \hdashline
 5 & 5+1 & 5+2 & 5+3 & 5+4 & 5+5 & 5+6 \\
 \hdashline
 6 & 6+1 & 6+2 & 6+3 & 6+4 & 6+5 & 6+6 \\
 \hdashline
\end{array}"

"1+3=2+2=3+1=4"

"2+6=3+5=4+4=5+3=6+2=8"

"6+6=12"

Let "X=" is the sum of the scores on the two dice


"P(X \text{is divided by 4})=\dfrac{9}{36}=\dfrac{1}{4}"